can someone help me with this question:
(f^o g)(x) = (g^o f)(x)
a. Always
b. Never
c. Sometimes
d. There is no such thing as (f^o g)(x)

Answer by Fombitz(32388)

Not sure what ^ means in (f^o g)(x) = (g^o f)(x).
I'd be inclined to answer d) no such thing.
.
.
.
If you mean composite functions and ^ is a typo,
f(g(x))=g(f(x)), then the answer is c) sometimes.